Circular No. 8374

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004df Further to IAUC 8370, T. Boles reports the discovery of an apparent

supernova (mag 18.4) on unfiltered CCD images taken on June 27.932 (at mag 18.4), 28.947 (mag 18.4), and July 13.001 (mag 18.1) UT. The new object is located at α = 15^{h}08^{m}20^{s}.65, δ = +21^{o}52′35′′.3, which is ≈ 10′′.7 west and 10′′.1 south of the center of its host galaxy. SN 2004df is not present on Boles’ images from May 18 and Mar. 20 (limiting mag 19.5), or on Digitized Sky Survey plates (red, 1993 Mar. 21; blue, 1986 Apr. 8).

SUPERNOVAE 2004by, 2004cz, AND 2004dd Further to IAUC 8372, R. J. Foley, H. Pugh, and A. V. Filippenko

(University of California) report that inspection of CCD spectra, obtained on July 18 UT, reveals that SN 2004by (IAUC 8349) is a peculiar type-II supernova. The object shows very narrow Hα emission (at most a few hundred km/s FWHM) and a corresponding low-velocity P-Cyg absorption line. SN 2004cz (IAUC 8368) and SN 2004dd (IAUC 8373) are normal type-II-P supernovae, with well-developed P-Cyg profiles of H, Ca II, Fe II, and other species. In SN 2004cz, the Na I D line (with a P-Cyg profile) is especially strong compared with that in SN 2004dd.

UNUSUAL CATACLYSMIC VARIABLE IN HERCULES A. Price, AAVSO, reports on observations of the unusual variable noted

on IAUC 8363 and 8369. A companion star < 1′′ away complicates measurements of its brightness; the two objects together were measured at V = 16 on July 16.217 UT, according to R. James, Las Cruces, NM. Further astrometry by A. Henden (1.0-m U.S. Naval Observatory telescope) gives the new variable’s position as α = 18^{h}39^{m}26^{s}.154, δ = +26^{o}04′09′′.89 (equinox 2000.0; ± 0′′.018); this is 1′′ north of its nearby companion, which has position end figures 26^{s}.195, 09′′.07 (identical within the errors to that of 2MASS 18392619+2604087 — itself having position end figures 26^{s}.19, 08′′.8; ± 0′′.06). A steady period of 0.0577 ± 0.0001 day exists during the outburst and decline. A search by Price and M. Templeton of 164 Harvard College Observatory plates taken from 1929 to 1950 reveals four previous outbursts on 1932 Apr. 21 (to blue mag ∼ 10.4), 1934 Oct. 17 (to mag ∼ 12.2), 1939 Aug. 16 (to mag ∼ 14.2, and 1941 Aug. 2 (to mag ∼ 10.6). Also, a nearby star (position end figures 27^{s}.37, 06′15′′.3) brightened by ∼ 0.2 mag/day from 2004 June 30.17 to July 9.24, when it was at V ∼ 14.

2004 July 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green